log7(3-4x)=log7(x/3)
Answer by stanbon(75887)

log7(3-4x)=log7(x/3)
-----------------
3-4x = x/3
4x + (x/3) = 3
12x + x = 9
13x = 9
x = 9/13
-------------
Cheers,
Stan H.